OpenStudy (anonymous):

Find all real solutions of the equation. (Enter your answers as a comma-separated list. If there is no real solution, enter NO REAL SOLUTION.) x4 + 6x3 + 6x2 = 0 @zbay

OpenStudy (anonymous):

Give us your thoughts, darling.

OpenStudy (anonymous):

So for this one it looks like you need to find a value for x that will make the left side equal the right side.

OpenStudy (anonymous):

yes i believe so

OpenStudy (anonymous):

Because you have a 4th order polynomial you will have 4 answers. However keep in mind that you may have multiple roots at the same value.

OpenStudy (anonymous):

So I would start by factoring out as much as you can and hopefully leaving you with a quadratic

OpenStudy (anonymous):

I'm terrible at factoring.

OpenStudy (anonymous):

Ok I will help you and hopefully you will see the process. When looking at the polynomial look at your lowest order term. This will hopefully allow you to identify the largest variable you can factor out.

OpenStudy (anonymous):

would x^4 be my lowest order term ?

OpenStudy (anonymous):

\[x^4+6x^3+6x^2=0\] We look at the last term and notice its x^2, and that we can factor that out of the rest of the terms. \[x^2(x^2+6x+6)=0\]

OpenStudy (anonymous):

okay i understand that part now

OpenStudy (anonymous):

X^4 would be your largest or highest order term. After factoring out x^2 you can then set both to zero \[x^2=0 \] \[x^2+6x+6=0\] at this point you can solve for x and this will give you the solutions for the problem

OpenStudy (anonymous):

I am guessing that you will need to use the quadratic equation for the 2nd term.

OpenStudy (anonymous):

i dont know what the second term is

OpenStudy (anonymous):

sorry I ment the 2nd equation \[x^2+6x+6=0\]

OpenStudy (anonymous):

are you talking about using the quadratic formula ?

OpenStudy (anonymous):

for the second equation

OpenStudy (anonymous):

Yes I would use the quadratic formula to solve for x on that equation. Or you can complete the square it looks like

OpenStudy (anonymous):

yeah ill use the quad. formula because i dont really remember how to do the completing the square

OpenStudy (anonymous):

I always tend to use the quadratic formula also

OpenStudy (anonymous):

remember your equation \[x^2=0\] will have 2 answers

OpenStudy (anonymous):

ok im working on it now

OpenStudy (anonymous):

ok great let me know if you have any problems

OpenStudy (anonymous):

my final answer is -6 (square root) 6 is that correct

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I'm getting x=-3-sqrt(3), and x=sqrt(3)-3

OpenStudy (anonymous):

Let me work it out for myself really quick

OpenStudy (anonymous):

I am going to upload a picture of my work and hopefully it will allow you to find your mistake. I'm guessing you forgot to divide by 2a

OpenStudy (anonymous):

okay i see what i did wrong , your right i didnt divide the whole equation by two

OpenStudy (anonymous):

awesome, did you have any other questions for me tonight?

OpenStudy (anonymous):

1. Find all real solutions of the equation. (Enter your answers as a comma-separated list. If there is no real solution, enter NO REAL SOLUTION.) x9 − 81x5 = 0 2.Find all real solutions of the equation. (Enter your answers as a comma-separated list. If there is no real solution, enter NO REAL SOLUTION.) ( square root) 6 − x + 1 = x − 3

OpenStudy (anonymous):

well this one is a little bit more tricky but we can apply the same idea

OpenStudy (anonymous):

1. x^9 -81x^5=0 ** 2.(sqaure root) 6-x + 1= x-3 **

OpenStudy (anonymous):

\[x^9-81x^5=0\] factor out x^5 \[x^5(x^4-81)=0\]

OpenStudy (anonymous):

So at this point we know we have 5 of the possible solutions with the term x^5=0 leaving us with 4 remaining solutions to find. \[x^4-81=0\]

OpenStudy (anonymous):

is the answer is 9 ??

OpenStudy (anonymous):

So now what do we get \[(x^2+9)(x^2-9)=0\]

OpenStudy (anonymous):

and we can factor out the term \[x^2-9=0\] to \[(x-3)(x+3)=0\]

OpenStudy (anonymous):

So now we just have to find the values of x for \[x^2+9=0\] I believe this will have an imaginary answer.

OpenStudy (anonymous):

imaginary number means it'll be something like 3i ?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

So the values for x are x=0,0,0,0,0,-3,3 then the imaginary roots

OpenStudy (anonymous):

so there are gonna be more than 3 answers ? what are all the zeros from ?

OpenStudy (anonymous):

Yes to find out how many answers there will be for a given polynomial look at the highest power in the polynomial.

OpenStudy (anonymous):

so for this one we have x^9 meaning there are nine given values of x that will satisfy the polynomial. Remember though some can repeat themselves for example when we factored out x^5=0 we know that the solution x=0 is repeated 5 times.

OpenStudy (anonymous):

(sqaure root) 6-x + 1= x-3 does it equal \[\sqrt{6-x+1}=x-3\]

OpenStudy (anonymous):

the sqaure root is just over 6-x

OpenStudy (anonymous):

\[\sqrt{6-x}+1=x-3\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Alright, move the 1 to the right hand side (RHS) isolating the radical on the left hand side (LHS). Square both sides eliminating the radical on the LHS giving you the term \[(x-4)^2\] that you will need to foil out. at this point in time you can move all terms to one side or the other and solve using the quadratic equation.

OpenStudy (anonymous):

im having trouble with that one , but i do have two more questions for you

OpenStudy (anonymous):

you there ?

OpenStudy (anonymous):

yes im still here

OpenStudy (anonymous):

1. Investments: if ben invest $4000 at 4% intrest per year how much additional money must he invest at 5 1/2 % annual interest to ensure that the intrest he recieves each year is 4 1/2 % ?

OpenStudy (anonymous):

2. Dimensions of a lot : A half acre bilding lot is five times as long as it is wide. What are its dimensions? ( 1 acre= 43,560ft^2)

OpenStudy (anonymous):

what interest formula have you been working with?

OpenStudy (anonymous):

I'm assuming the simple interest formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[I=PRT\]

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

1. Investments: if ben invest $4000 at 4% intrest per year how much additional money must he invest at 5 1/2 % annual interest to ensure that the intrest he recieves each year is 4 1/2 % ? ignore this just moving it down so I don't have to keep scorlling up

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Ok give me a min I'm working it out

OpenStudy (anonymous):

okay thankss

OpenStudy (anonymous):

Ok one down I will upload my work and you can look it over and ask any questions

OpenStudy (anonymous):

alright :)

OpenStudy (anonymous):

Dimensions of a lot : A half acre bilding lot is five times as long as it is wide. What are its dimensions? ( 1 acre= 43,560ft^2)

OpenStudy (anonymous):

your such a life saver. when i did it i was just dividing all types of numbers in the wrong spots .

OpenStudy (anonymous):

Yea the interest one throw me off for a min also

OpenStudy (anonymous):

I will be back in a few min if you have any questions

OpenStudy (anonymous):

okay thanks soo much

OpenStudy (anonymous):

not a problem do you understand how I worked out those last two?

OpenStudy (anonymous):

yes i see now , im more of a visual learner, if its just explained to me without working it out , i will not get it what so ever

OpenStudy (anonymous):

Ok great would you like for me to work out the problem with the square root in it? Didn't you say you were having problems with that one?

OpenStudy (anonymous):

\[\sqrt{6-x}+1=x-3\] \[\sqrt{6-x}=x-4\] \[(\sqrt{6-x})^2=(x-4)^2\] foil RHS, radical is eliminated \[6-x=x^2-8x+16\] \[0=x^2-7x+10\] \[0=(x-2)(x-5)\] \[x=2,5\]

OpenStudy (anonymous):

ehh im a little iffy about this problem, im confused but i dont really know what im confused out

OpenStudy (anonymous):

about*

OpenStudy (anonymous):

Well the only real trick to this problem is that radical, and whenever you square a radical it goes away leaving whats under it. After that part just group terms and factor with one of the methods you know.

OpenStudy (anonymous):

So are you good to go?